# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v1

@@ Appendix A Time Complexity
Appendix B Tau Recovery
Figure 7: Correlations of inferred rankings with planted rankings. (a) $\tau_{B}$ scores between the rankings inferred by the Bradley-Terry and partial rankings models with respect to the ground truth ranking for the case in which there are no partial rankings ($R=N=150$). (b) $\tau_{B}$ scores between the rankings inferred by the Bradley-Terry and partial rankings models with respect to the ground truth ranking for the case in which there are three planted partial rankings ($R=3$). All results are averaged over 10 different simulations and error bars represent 2 standard errors in the mean.
@@ Appendix B Tau Recovery
When partial rankings are present (panel (b)), the $\langle\tau_{B}\rangle$ value for the BT model plateaus around 0.8. This behaviour likely stems from the Bradley-Terry model’s inherent assumption of a complete ranking, which limits its ability to adapt once it has achieved the best possible alignment it can with the ground truth rankings. On the other hand, the PR model is able to obtain perfect correlation already at relatively low values of $\langle k\rangle$, thanks to its ability to account for and adapt to the presence of partial rankings in the data.
Appendix C Comparison with 1D Clustering

### v2

@@ Appendix B Time Complexity
Appendix C Tau Recovery
Figure 7: Correlations of inferred rankings with planted rankings. (a) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are no partial rankings ($R=N=150$). (b) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are three planted partial rankings ($R=3$). All results are averaged over 10 different simulations and error bars represent 2 standard errors in the mean.
@@ Appendix C Tau Recovery
When partial rankings are present (panel (b)), the $\langle\tau_{B}\rangle$ value for the BT model plateaus around 0.8. This behaviour likely stems from the BT model’s inherent assumption of a complete ranking, which limits its ability to adapt once it has achieved the best possible alignment it can with the ground truth rankings. On the other hand, the PR model is able to obtain perfect correlation already at relatively low values of $\langle k\rangle$, thanks to its ability to account for and adapt to the presence of partial rankings in the data.
Appendix D Comparison with 1D Clustering